OpenStudy (dj3strella):

i have a question 3. Which is a counterexample that disproves the conjecture? A student concludes that if x is a real number, then x ≥ x3. (Points : 1) x = 3 x = 1 x = 0 x = –1

OpenStudy (dj3strella):

@satellite73 can you help me please

hartnn (hartnn):

plug in your choices for x , one by one and see which one makes the inequality, x >= x^3 as FALSE. like if i take x=0 0>=0 is true so its not the correct answer. try same thing for other choices

OpenStudy (dj3strella):

so 3 > or = 3^3 1> or = 1^3 0 > or = 0^3 -1> or = -1^3

hartnn (hartnn):

which one of them is false ?

hartnn (hartnn):

3^3 = 27 -1^3 = -1

hartnn (hartnn):

is 3 greater than or = 27 ?

OpenStudy (dj3strella):

i dont remember @hartnn

hartnn (hartnn):

3 is lesser than 27 so 3 greater than or =27 is FALSE so x=3 is your counterexample :)

OpenStudy (dj3strella):

ok